OpenStudy (anonymous):

if p varies directly as q and p=18, q=36

OpenStudy (tkhunny):

"p varies directly as " MEANS p = k*q Just hear that in your head when you read it. Substitute the given values to find 'k'.

OpenStudy (anonymous):

thank you. this is my summer work for high school and I haven't been taught this yet.

OpenStudy (tkhunny):

You'll need this one, too. "p varies inversely as q" MEANS p = k/q

OpenStudy (anonymous):

thank you. I need help on this equation because I have a hard time with percentages. If a population of rats was 10 in January and 15 in February, what is the percent growth of these rats?

OpenStudy (tkhunny):

\(\dfrac{New-Old}{Old}\implies\dfrac{15-10}{10}\) Make that a percentage and you're done.